OpenStudy (dragons):

multiply and simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers ^3 sqrt y ^4 ^3 sqrt81y^5

OpenStudy (anonymous):

If you mean\[\sqrt[3]{y ^{4}}\times \sqrt[3]{81y ^{5}}\] then the working is \[y \sqrt[3]{y}\times 3y \sqrt[3]{y ^{2}}\] =\[3y ^{2}\times \sqrt[3]{y \times y ^{2}}\]=\[3y ^{2}timesy=3y ^{3}\]